You are standing at x = 8.0km. Lightning bolt 1 at x=0km and lightning bolt 2 at x = 13.0km. Both flashes reach your eye at the same time. Your assistant is standing at x = 2.0km If your assistant didn't see the flashes at the same time, what is the time difference between the two flashes?

answer in us
Answer #1

let \Delta t be the time difference between the two flashes; flash 1 flashes at time = t and flash 2 at time =t+ \Delta t

so the extra distance is traveled by flash 1 in  \Delta t time

i.e \Delta t=\frac{3000}{3*10^{8}}=10 miliseconds

for assistant time gap =\frac{9000}{3*10^{8}}+\Delta t=30 +10=40miliseconds